When electing an office bearer in a company the votes are normally distributed with mean μ and standard deviation σ. More than 85 people voting is 5%. Less than 27 people voting is 15%. Find the mean μ and standard deviation σ.


1
Expert's answer
2021-09-30T09:29:15-0400

"z=\frac{x-\mu}{\sigma}"


"P(x>85)=1-P(z<z_1=\frac{85-\mu}{\sigma})=0.05"

"P(x<27)=P(z<z_2=\frac{27-\mu}{\sigma})=0.15"


From z-table:

"z_1=1.645"

"z_2=-2.17"


"\frac{85-\mu}{\sigma}=1.645"

"\frac{27-\mu}{\sigma}=-2.17"


"\mu=85-1.645\sigma"

"27-85+1.645\sigma=-2.17\sigma"

"3.81\sigma=58"


"\sigma=58/3.81=15.22"

"\mu=85-1.645\cdot 15.22=60"